### v2

# Mixed Berndt-Type Integrals and Generalized Barnes Multiple Zeta Functions

Abstract. In this paper, we define and study four families of Berndt-type integrals, called mixed Berndt-type integrals, which contains (hyperbolic) sine and cosine functions in the integrand function. By contour integration, these integrals are first converted to some hyperbolic (infinite) sums of Ramanujan type, all of which can be calculated in closed forms by comparing both the Fourier series expansions and the Maclaurin series expansions of a few Jacobi elliptic functions. These sums can be expressed as rational polynomials of $\Gamma(1/4)$ and $\pi^{-1}$ which give rise to the closed formulas of the mixed Berndt-type integrals we are interested in. Moreover, we also present some interesting consequences and illustrative examples. Further, we define a generalized Barnes multiple zeta function, and find a classic integral representation of the generalized Barnes multiple zeta function. Furthermore, we give an alternative evaluation of the mixed Berndt-type integrals in terms of the generalized Barnes multiple zeta function. Finally, we obtain some direct evaluations of rational linear combinations of the generalized Barnes multiple zeta function.
Keywords: Berndt-type integral, Hyperbolic and trigonometric functions, Contour integration, Jacobi elliptic functions, Generalized Barnes multiple zeta functions.
AMS Subject Classifications (2020): 33E05, 33E20, 44A05, 11M99.
1 Introduction
Let $m>0$ be an integer. The Berndt-type integrals of order $p$ are defined by the form (see Bradshaw-Vignat [7] and Xu-Zhao [22])
$I_{\pm}(s,m):=\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{s-1}\mathrm{d}x}{(\cos x\pm\cosh x)^{m}},$ (1.1)
where $s\geq 1$ and $m\geq 1$ if the denominator has “$+$’ sign, and $s\geq 2m+1$ otherwise. This kind of improper integral can be traced back to the esteemed mathematician S. Ramanujan. First as a problem submitted by Ramanujan [14, pp. 325-326] over a century ago to the Indian Journal of Pure and Applied Mathematics:
$\int_{0}^{\infty}\displaystyle\frac{\sin(nx)}{x(\cos x+\cosh x)}\mathrm{d}x=\displaystyle\frac{\pi}{4}.$
Here $n$ is any odd positive integer. Four year after submission, Wilkinson [19] provided a proof. About 80 years later, Ismail and Valent [10] found the mystery integral (Ismail-type integrals)
$\displaystyle\int_{-\infty}^{\infty}{\displaystyle\frac{\mathrm{d}t}{\cos\left(K\sqrt{t}\right)+\cosh\left(K^{\prime}\sqrt{t}\right)}=2},$ (1.2)
where $K\equiv K(x)$ denotes the complete elliptic integral of the first kind and $K^{\prime}=K(1-x)$, see Section 2. The integrand in (1.2) is the weight function for a sequence of polynomials arising in the Hamburger moment problem studied by Ismail and Valent. The systematic study of Berndt-type integrals began with the works of Kuznetsov [12] and Berndt [6] about ten years ago. Some recent results on infinite series involving Berndt-type integrals may be found in the works of [7, 13, 16, 17, 21, 22, 24] and the references therein. Kuznetsov [12] used the contour integrations and theta functions to prove a more general result of Ismail-type integrals
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{t^{n}\mathrm{d}t}{\cos(K\sqrt{t})+\cosh(K^{\prime}\sqrt{t})}=(-1)^{n}\displaystyle\frac{\mathrm{d}^{2n+1}}{\mathrm{d}u^{2n+1}}\displaystyle\frac{{\rm sn}(u,k)}{{\rm cd}(u,k)}|_{u=0},$ (1.3)
where $x=k^{2}\ (0<k<1)$ and the Jacobi elliptic function ${\rm sn}(u,k)$ is defined via the inversion of the elliptic integral
$u=\displaystyle\int_{0}^{\varphi}\displaystyle\frac{\mathrm{d}t}{\sqrt{1-k^{2}\sin^{2}t}}\quad(0<k^{2}<1),$ (1.4)
namely, ${\rm sn}(u):=\sin\varphi$. As before, we refer $k$ as the elliptic modulus. We also write $\varphi={\rm am}(u,k)={\rm am}(u)$ and call it the Jacobi amplitude. The Jacobi elliptic function ${\rm cd}(u,k)$ may be defined as follows:
${\rm cd}(u,k):=\displaystyle\frac{{\rm cn}(u,k)}{{\rm dn}(u,k)},$
where ${\rm cn}(u,k):=\sqrt{1-{\rm sn}^{2}(u,k)}$ and ${\rm dn}(u,k):=\sqrt{1-k^{2}{\rm sn}^{2}(u,k)}$. Berndt [6] provided a direct evaluation of (1.2) and other similar integrals of this type by using Cauchy’s residue theory, the Fourier series expansions, and the Maclaurin series expansions of a few Jacobi elliptic functions. In particular, using Berndt’s results [6] and Xu-Zhao’s result [22], it is easy to see that the Berndt-type integrals of order one have the following structures
$I_{-}(4p,1)\in\mathbb{Q}\displaystyle\frac{\Gamma^{8p}(1/4)}{\pi^{2p}}\quad(p\in\mathbb{N})\quad\text{and}\quad I_{+}(4p+2,1)\in\mathbb{Q}\displaystyle\frac{\Gamma^{8p+4}(1/4)}{\pi^{2p+1}}\quad(p\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}).$
Moreover, Berndt [6, Thm. 6.1] also considered some integrals contain $\sinh x\pm i\sin x$ in their integrands instead of $\cosh x\pm\cos x$. Further, using the contour integrals and the Fourier series expansion and Maclaurin series expansion of a certain Jacobi elliptic function, Rui-Xu-Yang-Zhao obtained a structural result on the original Berndt type integral involving (hyperbolic) sine functions (see [16, Thm. 3.4])
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{2m}(\sinh x-(-1)^{m}\sin x)}{\sinh^{2}x+\sin^{2}x}\mathrm{d}x\in\mathbb{Q}\displaystyle\frac{\Gamma^{4m+2}(1/4)}{\pi^{(2m+1)/2}}\sqrt{2}\quad(m\in\mathbb{N})$ (1.5)
and prove the following result involving another family of Berndt-type integrals (see [16, Cor. 6.5]):
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{4p+1}(\sin^{2}x-\sinh^{2}x)}{\left(\sinh^{2}x+\sin^{2}x\right)^{2}}\mathrm{d}x\in\mathbb{Q}\displaystyle\frac{\Gamma^{8p+4}(1/4)}{\pi^{2p+2}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8p+8}(1/4)}{\pi^{2p+4}}\quad(p\in\mathbb{N}).$ (1.6)
Recently, by extending an argument as used in the proof of the main theorems of [6], Xu and Zhao [21] proved that the following explicit evaluations of two families of Berndt-type integrals of order two (cf. [21, Thm. 1.3]):
$I_{\pm}(4p+2,2)\in\mathbb{Q}\displaystyle\frac{\Gamma^{8p}(1/4)}{\pi^{2p}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8p+8}(1/4)}{\pi^{2p+4}}\quad(p\in\mathbb{N}).$
Further, Rui, Xu and Zhao [17, Thm. 1.1] established explicit evaluations of two families of Berndt-type integrals of order three by special values of the Gamma function. More generally, let $X=\Gamma^{4}(1/4)$ and $Y=\pi^{-1}$, and $\mathbb{Q}[x,y]$ denotes all rational linear combinations of products of $x$ and $y$ (the polynomial ring generated by $x$ and $y$ with rational coefficients). Xu and Zhao [21, 22] proved two structural theorems on the general Berndt-type integrals with the denominator having arbitrary positive degrees. Specifically, they shown that for all integers $m\geq 1$ and $p\geq[m/2]$, the Berndt-type integrals satisfy
$I_{+}(4p+2,m)\in\mathbb{Q}[X,Y],$
where the degrees of $X$ have the same parity as that of $m$ and are between $22p-m+$ 2 and $2p+m$, inclusive, while the degrees of $Y$ are between $22p-m+$ 2 and $22p+3m-$ 2, inclusive. Similarly, they shown that
$I_{-}(4p,2m+1)\in\mathbb{Q}[X,Y]$
for all $p\geq m+1\geq 1$ and that the degrees of $X$ in each term are always even and between $2p-2m$ and $2p+2m$, while the degrees of $Y$ are between $2p-2m$ and $2p+6m$. In the even order case, they shown that
$I_{-}(4p+1,2m)\in\mathbb{Q}[X,Y]$
for all integers $p\geq m\geq 1$ and the powers of $X$ are between $2p+2-2m$ and $2p+2m$, while the degrees of $Y$ are between $2p+2-2m$ and $22p+6m-$ 2. Very recently, it is very surprising that Bradshaw and Vignat [7, Cor. 2] discovered a close connection between Berndt-type integrals and Barnes zeta functions. Moreover, Bradshaw and Vignat [7, Cor. 5] used the Kuznetsov’s method to prove the following relations between Jacobi elliptic functions and another Ismail-type integrals: for $n\in\mathbb{N}_{0}$,
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{x^{n+1}dx}{\cos(\sqrt{x}K)-\cosh(\sqrt{x}K^{\prime})}=(-1)^{n+1}4\displaystyle\frac{d^{2n+1}}{du^{2n+1}}\displaystyle\frac{{\rm sn}^{2}(u,k)}{{\rm cd}^{2}(u,k){\rm sd}(2u,k)}|_{u=0},$ (1.7)
where the Jacobi elliptic function ${\rm sd}(u)\equiv{\rm sd}(u,k)$ defined by
${\rm sd}(u,k):=\displaystyle\frac{{\rm sn}(u,k)}{{\rm dn}(u,k)}.$ (1.8)
From the above, it can be seen that the integrand in the Berndt-type integrals studied above either only contain (hyperbolic) sine functions or (hyperbolic) cosine functions. In this paper, we consider the following infinite integrals containing (hyperbolic) sine functions and (hyperbolic) cosine functions in their integrands:
$B^{\pm}_{\pm}(s):=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{s}\left(\sinh x\pm\sin x\right)\mathrm{d}x}{\left(\sinh^{2}x+\sin^{2}x\right)\left(\cosh x\pm\cos x\right)}}.$ (1.9)
We call them mixed Berndt-type integrals. We use the contour integrals and the series associated with Jacobi elliptic functions to establish two structural theorems of mixed Berndt-type integrals. We will prove the following evaluations: let $\Gamma:=\Gamma(1/4)$, for $m\in\mathbb{N}$,
$B^{-}_{+}(4m)=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left(\sinh^{2}x+\sin^{2}x\right)\left(\cosh x+\cos x\right)}}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$ (1.10)
$B^{+}_{+}(4m-2)=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sin x+\sinh x\right)\mathrm{d}x}{\left(\sinh^{2}x+\sin^{2}x\right)\left(\cosh x+\cos x\right)}}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}},$ (1.11)
and
$B^{+}_{-}(4m)=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left(\sinh^{2}x+\sin^{2}x\right)\left(\cosh x-\cos x\right)}}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$ (1.12)
$B^{-}_{-}(4m-2)=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left(\sinh^{2}x+\sin^{2}x\right)\left(\cosh x-\cos x\right)}}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}}\quad(m\geq 2).$ (1.13)
Furthermore, we define the generalized Barnes multiple zeta function by:
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N}):=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\displaystyle\frac{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}},}\quad(\Re(\omega)>\Re(s)>N),$
where $\boldsymbol{\sigma}_{N}=\left(\sigma_{1},\ldots,\sigma_{N}\right)\in\left\{\pm 1\right\}^{N}$. We find a classic integral representation:
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})=\displaystyle\frac{1}{\Gamma\left(s\right)}\displaystyle\int_{0}^{\infty}{u^{s-1}}e^{-wu}\prod_{j=1}^{N}{(}1-\sigma_{j}e^{-a_{j}u})^{-1}\mathrm{d}u,$
and establish some explicit relations between the mixed Berndt-type integrals and (generalized) Barnes multiple zeta functions. Finally, we establish some closed form evaluations of the combinations of several generalized Barnes multiple zeta function.
2 Some Preliminary Results
Let ${}_{2}F_{1}\left(a,b;c;x\right)$ denote the Gaussian hypergeometric function that
${{}_{2}}F_{1}(a,b;c;x)=\displaystyle\sum\limits_{n=0}^{\infty}\displaystyle\frac{(a)_{n}(b)_{n}}{(c)_{n}}\displaystyle\frac{x^{n}}{n!}\quad(a,b,c\in\mathbb{C}).$
Here $\left(a\right)_{0}:=1$ and $\left(a\right)_{n}=\displaystyle\frac{\Gamma\left(a+n\right)}{\Gamma\left(a\right)}=\prod_{j=0}^{n-1}{\left(a+j\right)}\quad(n\in\mathbb{N}),$ and $\Gamma(x)$ is the Gamma function defined by
$\Gamma\left(z\right):=\displaystyle\int_{0}^{\infty}{e^{-t}t^{z-1}\mathrm{d}t}\quad\left(\Re\left(z\right)>0\right).$
We know that ${}_{2}F_{1}\left(a,b;c;x\right)$ has numerous applications in mathematics, physics and even engineering. In particular, it can be used to evaluate some important elliptic integrals. For example, the complete elliptic integral of the first kind is defined by
$K:=K\left(k^{2}\right)=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right),$
where $x=k^{2}\quad(0<k<1)$ is the modulus of $K$. Let $k^{\prime}=\sqrt{1-k^{2}}$, and
$K^{\prime}:=K\left(k^{\prime 2}\right)=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{\prime 2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;1-k^{2}\right).$
Similarly, the complete elliptic integral of the second kind is denoted by (Whittaker and Watson [18])
$E:=E(k^{2}):=\displaystyle\int_{0}^{\pi/2}\sqrt{1-k^{2}\sin^{2}\varphi}\mathrm{d}\varphi=\displaystyle\frac{\pi}{2}{{}_{2}}F_{1}\left(-\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right).$
Adopting Ramanujan’s notations, we let
$x:=k^{2},\ y\left(x\right):=\pi K^{\prime}/K,\ q\equiv q\left(x\right):=e^{-y},\ z:=z\left(x\right)=2K/\pi,\ z^{\prime}=dz/dx.$ (2.14)
Then
$\frac{\mathrm{d}^{n}z}{\mathrm{d}x^{n}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}}{n!}\,_{2}F_{1}\left(\displaystyle\frac{1}{2}+n,\displaystyle\frac{1}{2}+n,1+n;x\right).$
Applying the identity ([2])
$\,{}_{2}F_{1}\left(a,b,\displaystyle\frac{a+b+1}{2};\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma\left(\displaystyle\frac{1}{2}\right)\Gamma\left(\displaystyle\frac{a+b+1}{2}\right)}{\Gamma\left(\displaystyle\frac{a+1}{2}\right)\Gamma\left(\displaystyle\frac{b+1}{2}\right)},$
we can calculate that
$\frac{d^{n}z}{dx^{n}}\mid_{x=\tfrac{1}{2}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}\sqrt{\pi}}{\Gamma^{2}\left(\displaystyle\frac{n}{2}+\displaystyle\frac{3}{4}\right)}.$
3 Some Definitions and Lemmas
Definition 3.1.
Let $m\in\mathbb{N}$ and $p\in\mathbb{Z}$. Define
$S_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{n^{p}}{\sinh^{m}(ny)},$ $\overline{S}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{n^{p}}{\sinh^{m}(ny)}(-1)^{n-1},$
$S_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(2n-1)^{p}}{\sinh^{m}((2n-1)y/2)},$ $\overline{C}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{n^{p}}{\cosh^{m}(ny)}(-1)^{n-1},$
$\mathrm{DS}_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(2n-1)^{p}\cosh((2n-1)y/2)}{\sinh^{m}((2n-1)y/2))}},$ $\mathrm{DS}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{n^{p}\cosh\left(ny\right)}{\sinh^{m}(ny)}}.$
Lemma 3.1.
([21, Lem. 7.3.]) For any integer $m\geq 0$ there is a polynomial $g_{m}(X,Y,Z)\in\mathbb{Z}[X,Y,Z]$ such that
$\displaystyle\frac{\mathrm{d}^{m}}{\mathrm{d}u^{m}}\mathrm{sn}\left(u\right)=g_{m}\left(\sin\varphi,\cos\varphi,k^{2}\right)\cdot\begin{cases}\left(1-k^{2}\sin^{2}\varphi\right)^{1/2},if\,\,m\,\,is\,\,odd;\\ 1,if\,\,m\,\,is\,\,even,\\ \end{cases}$ (3.15)
Lemma 3.2.
([23, Thm.2.4]) Let $x,y,z$ and $z^{\prime}$ satisfy (2.14). Given the formula $\Omega(x,e^{-y},z,z^{\prime})=0$, we have the transformation formulas
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0,$ (3.16)
$\Omega\left(\displaystyle\frac{x}{x-1},-e^{-y},z\sqrt{1-x},\left(1-x\right)^{\tfrac{3}{2}}\left(\displaystyle\frac{z}{2}-\left(1-x\right)z^{\prime}\right)\right)=0,$ (3.17)
$\Omega\left(\left(\displaystyle\frac{1-\sqrt{1-x}}{1+\sqrt{1-x}}\right)^{2},e^{-2y},\displaystyle\frac{z\left(1+\sqrt{1-x}\right)}{2},\displaystyle\frac{\left(1+\sqrt{1-x}\right)^{3}}{4\left(1-\sqrt{1-x}\right)}\left(\left(1-x+\sqrt{1-x}\right)z^{\prime}-\displaystyle\frac{z}{2}\right)\right)=0.$ (3.18)
Lemma 3.3.
([21]) Let $n\in\mathbb{Z}$. We have
$\displaystyle\frac{\left(-1\right)^{n}}{\sin\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1\mp i\right)n\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)n\pi}+2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)n\pi\right)^{2k-1}},$ (3.19)
$\displaystyle\frac{\left(-1\right)^{n-1}}{\cos\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1-i\right)\tilde{n}\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)\tilde{n}\pi}-2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)\tilde{n}\pi\right)^{2k-1}},$ (3.20)
where $\tilde{n}:=n-\displaystyle\frac{1}{2}$. Here $\zeta(s)$ and $\overline{\zeta}(s)$ stand for the Riemann zeta function and alternating Riemann zeta function, respectively, which are defined by
$\zeta\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{1}{n^{s}}}\quad(\Re(s)>1)\quad\text{and}\quad\bar{\zeta}\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n-1}}{n^{s}}}\quad\left(\Re\left(s\right)>0\right).$
When $s=2m$ is an even, Euler proved the famous formula
$\zeta\left(2m\right)=\displaystyle\frac{\left(-1\right)^{m-1}B_{2m}\left(2\pi\right)^{2m}}{2\left(2m\right)!},$
where $B_{2m}\in\mathbb{Q}$ are Bernoulli numbers defined by the generating function
$\frac{x}{e^{x}-1}=\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{B_{n}}{n!}x^{n}.}$
Lemma 3.4.
([23, Lem. 1]) Let $p$ be a positive integer and $\alpha,\beta$ be real numbers such as $\alpha\beta=\pi^{2}$
$\alpha^{2p+1}\bar{C}_{2p,2}\left(\alpha\right)-\left(-1\right)^{p}p\displaystyle\frac{\pi^{2p}}{2^{2p-2}}S^{\prime}_{2p-1,1}\left(\beta\right)+\left(-1\right)^{p}\displaystyle\frac{\pi^{2p}}{2^{2p}}\beta\mathrm{DS}^{\prime}_{2p,2}\left(\beta\right)=0,$ (3.21)
$\alpha^{2p}\bar{S}_{2p,2}\left(\alpha\right)-2p\left(-1\right)^{p-1}\pi^{2p-2}\beta S_{2p-1,1}\left(\beta\right)-\left(-1\right)^{p}\pi^{2p-2}\beta^{2}\mathrm{DS}_{2p,2}\left(\beta\right)-\delta_{p}=0,$ (3.22)
where $\delta_{1}=\displaystyle\frac{1}{2}$ and $\delta_{p}=0$ if $p\geq 2$.
Lemma 3.5.
([1, Thm.2]) The Maclaurin series of ${\rm sn}(u)$ has the form
$\mathrm{sn}\left(u\right)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{q}_{2n+1}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n+1}}{\left(2n+1\right)!}},$ (3.23)
where $\text{q}_{2n+1}(x)=\displaystyle\sum^{n}_{j=0}a_{j}x^{j}$ and $a_{j}=a_{n-j}\ (j=0,\ldots,n$).
4 Mixed Berndt-Type Integrals via Hyperbolic Series
In this section, we will establish some explicit relations between the mixed Berndt-type integrals and some hyperbolic summations by using the contour integrations.
Theorem 4.1.
For any $a\geq 1$, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $-i^{a}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=\left(\displaystyle\frac{1+i}{2}\pi\right)^{a+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}n^{a}}{\cosh^{2}\left[\tfrac{n\pi}{2}\right]}}.$ (4.24)
Proof.
Let $z=x+iy$ for $x,y\in\mathbb{R}$. Consider
$\mathscr{I}_{a}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z+\cos z\right)}}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{F_{1}\left(z\right)}\mathrm{d}z,$ (4.25)
where $C_{R}$ denotes the quarter- circular contour consisting of the interval $[0,R]$, the qiarter-circle $\Gamma_{R}$ with $|z|=R,0\leq argz\leq\pi/2$, and $[iR,0]$, please see the following figure:
Obviously, there exist poles in case that
$\left(\sinh z+i\sin z\right)\left(\cosh x+\cos x\right)=2i\cos^{2}\left\{\displaystyle\frac{1+i}{2}z\right\}\sin\left\{\left(1-i\right)z\right\}=0.$
The only lying inside $C_{R}$ are $z_{n}=n(1+i)\pi/2\ (n\geq 1,\ |z_{n}|<R)$. Using Lemma 3.3, the residue $\text{Res}[F_{1}(z),z_{n}]$ at each such pole is give by
$\mathrm{Res}\left[F\left(z\right),z_{n}\right]=\displaystyle\frac{\left(-1\right)^{n}\left[\tfrac{1+i}{2}n\pi\right]^{a}}{2i\left(1-i\right)\cos^{2}\left[\left(\tfrac{1+i}{2}\right)^{2}n\pi\right]}$
$=\left(\displaystyle\frac{1+i}{2}\right)^{a+1}\displaystyle\frac{\pi^{a}}{2i}\displaystyle\frac{\left(-1\right)^{n}n^{a}}{\cos^{2}\left[\tfrac{i}{2}n\pi\right]}=\left(\displaystyle\frac{1+i}{2}\right)^{a+1}\displaystyle\frac{\pi^{a}}{2i}\displaystyle\frac{\left(-1\right)^{n}n^{a}}{\cosh^{2}\left[\tfrac{n\pi}{2}\right]}.$ (4.26)
As $R\rightarrow\infty$, it is to show that
$\int_{\Gamma_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z+\cos z\right)}}=o\left(1\right).$
Applying the residue theorem and letting $R\rightarrow\infty$, we conclude that
$2\pi i\displaystyle\sum_{n=1}^{\infty}{\mathrm{Res}\left[F_{1}\left(z\right),z_{n}\right]}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z+\cos z\right)}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sinh x+i\sin x\right]\left[\cosh x+\cos x\right]}}+i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(ix\right)^{a}\mathrm{d}x}{\left[\sinh\left(ix\right)+i\sin\left(ix\right)\right]\left[\cosh\left(ix\right)+\cos\left(ix\right)\right]}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}-i^{a}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sin x+i\sinh x\right]\left[\cosh x+\cos x\right]}}.$
Thus, combining related identities, we deduce the desired result. ∎
Theorem 4.2.
For any $a\geq 3$, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $\left(-1\right)^{a}i^{a}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=\left(\displaystyle\frac{1-i}{2}\pi\right)^{a+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}n^{a}}{\sinh^{2}\left(\displaystyle\frac{n\pi}{2}\right)}}.$ (4.27)
Proof.
Consider
$\mathscr{J}_{a}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)}}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{F_{2}\left(z\right)}\mathrm{d}z,$ (4.28)
where $C_{R}$ denotes the quarter-circular contour consisting of the interval $[0,R]$, the qiarter-circle $\Gamma_{R}$ with $|z|=R\ (-\pi/2\leq argz\leq 0)$, and $[-iR,0]$, please see the following figure.
Obviously, there exist poles in case that
$\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)=-2i\sin^{2}\left\{\displaystyle\frac{1-i}{2}z\right\}\sin\left\{\left(1+i\right)z\right\}=0.$
The only lying inside $C_{R}$ are $z_{n}=n(1-i)\pi/2\ (n\geq 1,\ |z_{n}|<R)$. Using Lemma 3.3, the residue $\text{Res}[F_{2}(z),z_{n}]$ at each such pole is give by
$\mathrm{Res}\left[F_{2}\left(z\right),z=\displaystyle\frac{1-i}{2}n\pi\right]=\displaystyle\frac{\left(-1\right)^{n}\left(\tfrac{1-i}{2}n\pi\right)^{a}}{2i\left(1+i\right)\sinh^{2}\left(\left(\tfrac{1-i}{2}\right)^{2}n\pi\right)}$
$=\displaystyle\frac{\left(-1\right)^{n}\left(\tfrac{1-i}{2}n\pi\right)^{a}}{2i\left(1+i\right)\sinh^{2}\left(\tfrac{n\pi}{2}\right)}=\left(\displaystyle\frac{1-i}{2}\right)^{a+1}\displaystyle\frac{\pi^{a}}{2i}\displaystyle\frac{\left(-1\right)^{n}n^{a}}{\sinh^{2}\left(\displaystyle\frac{n\pi}{2}\right)}.$ (4.29)
It is to show that, letting $R\rightarrow\infty$,
$\int_{\Gamma_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)}}=o\left(1\right).$
Applying the residue theorem and letting $R$ tend to infinity, we conclude that
$2\pi i\displaystyle\sum_{n=1}^{\infty}{\mathrm{Res}\left[F_{2}\left(z\right),z_{n}\right]}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sinh x+i\sin x\right]\left[\cosh x-\cos x\right]}}-i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(-ix\right)^{a}\mathrm{d}x}{\left[\sinh\left(ix\right)+i\sin\left(ix\right)\right]\left[\cosh\left(ix\right)-\cos\left(ix\right)\right]}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+\displaystyle\int_{0}^{\infty}{\displaystyle\frac{(-i)^{a}x^{a}\mathrm{d}x}{\left[\sin x+i\sinh x\right]\left[\cosh x-\cos x\right]}}.$
Thus, the proof is completed. ∎
Theorem 4.3.
For any $a\geq 2$, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sinh x-i\sin x\right]\left[\cosh x-\cos x\right]}}-\left(-i\right)^{a+1}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sinh x+i\sin x\right]\left[\cosh x-\cos x\right]}}$
$=2\left(1-i\right)^{a-1}\pi^{a}\displaystyle\sum_{n=1}^{\infty}{\left[\displaystyle\frac{an^{a-1}}{\sinh\left(2n\pi\right)}-2\pi\displaystyle\frac{n^{a}\cosh\left(2n\pi\right)}{\sinh^{2}\left(2n\pi\right)}\right]}.$ (4.30)
Proof.
Consider
$\mathscr{H}_{a}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z-i\sin z\right)\left(\cosh z-\cos z\right)}},$ (4.31)
where $C_{R}$ denotes the same positively oriented quarter-circular contour that we used in the proof of Theorem 4.2. The proof is completely similar as the proof of Theorem 4.2. So we leave it to the interested reader. ∎
Theorem 4.4.
For any $a\geq 1$, we have
$\displaystyle\int_{0}^{\infty}$ ${\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sinh x-i\sin x\right]\left[\cosh x+\cos x\right]}}-\displaystyle\int_{0}^{\infty}{\displaystyle\frac{i^{a+1}x^{a}\mathrm{d}x}{\left[\sinh x+i\sin x\right]\left[\cosh x+\cos x\right]}}$
$\quad\quad=\displaystyle\frac{\left(1+i\right)^{1+a}i}{2^{a-1}}\pi^{a}\displaystyle\sum_{n=0}^{\infty}{\left[\displaystyle\frac{a\left(2n+1\right)^{a-1}}{\sinh\left(\left(2n+1\right)\pi\right)}-\pi\displaystyle\frac{\left(2n+1\right)^{a}\cosh\left(\left(2n+1\right)\pi\right)}{\sinh^{2}\left(\left(2n+1\right)\pi\right)}\right]}.$ (4.32)
Proof.
Consider
$\mathscr{W}_{a}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z-i\sin z\right)\left(\cosh z+\cos z\right)}},$ (4.33)
where $C_{R}$ denotes the same positively oriented quarter-circular contour that we used in the proof of Theorem 4.1. The proof is completely similar as the proof of Theorem 4.1. We leave the details to the interested reader. ∎
Setting $a=2\mu+1$ and $2\mu$ in Theorems 4.1-4.4 respectively, yield the following corollaries.
Corollary 4.5.
For $\mu\in\mathbb{N}$, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left[1+\left(-1\right)^{\mu+1}\right]x^{2\mu+1}\sinh x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}+$ $i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left[\left(-1\right)^{\mu+1}-1\right]x^{2\mu+1}\sin x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=-\left(\displaystyle\frac{i}{2}\pi^{2}\right)^{\mu+1}\bar{C}_{2\mu+1,2}\left(\displaystyle\frac{\pi}{2}\right),$ (4.34)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(1+\left(-1\right)^{\mu+1}\right)x^{2\mu+1}\sinh x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(\left(-1\right)^{\mu+1}-1\right)x^{2\mu+1}\sin x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=-\left(-\displaystyle\frac{i}{2}\pi^{2}\right)^{\mu+1}\bar{S}_{2\mu+1,2}\left(\displaystyle\frac{\pi}{2}\right),$ (4.35)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(1+\left(-1\right)^{\mu}\right)x^{2\mu+1}\sinh x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}+}i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(1-\left(-1\right)^{\mu}\right)x^{2\mu+1}\sin x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{i^{\mu+2}\pi^{2\mu+1}}{2^{\mu-1}}\left[\left(2\mu+1\right)S^{\prime}_{2\mu,1}\left(2\pi\right)-\pi\mathrm{DS}^{\prime}_{2\mu+1,2}\left(2\pi\right)\right],$ (4.36)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(1+\left(-1\right)^{\mu}\right)x^{2\mu+1}\sinh x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(1+\left(-1\right)^{\mu+1}\right)x^{2\mu+1}\sin x\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=\left(-i\right)^{\mu}2^{\mu+1}\pi^{2\mu+1}{\left[\left(2\mu+1\right)S_{2\mu,1}\left(2\pi\right)-2\pi\mathrm{DS}_{2\mu+1,2}\left(2\pi\right)\right]}.$ (4.37)
Corollary 4.6.
For $\mu\in\mathbb{N}$, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}-$ $\left(-1\right)^{\mu}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=-\left(\displaystyle\frac{1+i}{2}\pi\right)^{2\mu+1}\bar{C}_{2\mu,2}\left(\displaystyle\frac{\pi}{2}\right),$ (4.38)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $\left(-1\right)^{\mu}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=-\left(\displaystyle\frac{1-i}{2}\pi\right)^{2\mu+1}\bar{S}_{2\mu,2}\left(\displaystyle\frac{\pi}{2}\right),$ (4.39)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x+i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}-\left(-1\right)^{\mu}i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$\quad\quad=\displaystyle\frac{\left(1+i\right)^{2\mu+1}i}{2^{2\mu-1}}\pi^{2\mu}\left[2\mu S^{\prime}_{2\mu-1,1}\left(2\pi\right)-\pi\mathrm{DS}^{\prime}_{2\mu,2}\left(2\pi\right)\right],$ (4.40)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x+i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+\left(-1\right)^{\mu}i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=4\left(1-i\right)^{2\mu-1}\pi^{2\mu}\left[\mu S_{2\mu-1,1}\left(2\pi\right)-\pi\mathrm{DS}_{2\mu,2}\left(2\pi\right)\right].$ (4.41)
5 Evaluations of Hyperbolic Summations via Jacobi Functions
Now, we evaluate the the reciprocal hyperbolic series of Ramanujan type
$\bar{C}_{2p,2}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{n^{2p}\left(-1\right)^{n-1}}{\cosh^{2}\left(ny\right)}}\quad\text{and}\quad\bar{S}_{2p,2}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{n^{2p}\left(-1\right)^{n-1}}{\sinh^{2}\left(ny\right)}}$
by applying the Fourier series expansions and the Maclaurin series expansions of a few Jacobi elliptic functions. Let $x,y$ and $z$ satisfy the relations in (2.14).
Theorem 5.1.
Let $p\in\mathbb{N}$, we have
$\bar{C}_{2p,2}\left(y\right)=$ $\left(-1\right)^{p}\displaystyle\frac{pz^{2p+1}z^{\prime}x(1-x)}{2^{2p-1}}\mathrm{q}_{2p-1}\left(1-x\right)\sqrt{1-x}$
$+\left(-1\right)^{p}\displaystyle\frac{z^{2p+2}x(1-x)}{2^{2p}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\mathrm{q}_{2p-1}\left(1-x\right)\sqrt{1-x}\right],$ (5.42)
where ${\rm q}_{2p-1}\left(x\right)$ appears in the coefficients of the Maclaurin series of $\mathrm{sn}\left(u\right)$.
Proof.
Setting $\alpha=y$ and $\beta=\tfrac{\pi^{2}}{y}$ in Lemma 3.4 gives
$\bar{C}_{2p,2}\left(y\right)=\displaystyle\frac{\left(-1\right)^{p}}{y^{2p+1}}\displaystyle\frac{p\pi^{2p}}{2^{2p-2}}S_{2p-1,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)-\left(-1\right)^{p}\displaystyle\frac{\pi^{2p+2}}{2^{2p}y^{2p+2}}\text{DS}^{\prime}_{2p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.43)
Now, we need to compute the series $S_{2p-1,1}^{{}^{\prime}}\left(y\right)$. According to Lemma 3.1 or Lemma LABEL:sn_coefficient_symmetry, there are polynomials where ${\rm q}_{2n-1}(x)\in\mathbb{Z}[x]$ such that the Maclaurin series of ${\rm sn}(u)$ has the form
$\mathrm{sn}\left(u\right)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{q}_{2n+1}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n+1}}{\left(2n+1\right)!}}.$ (5.44)
From [8], we record the Fourier series expansions of Jacobi elliptic functions ${\rm sn}(u)$
${\rm sn}(u)=\displaystyle\frac{2\pi}{Kk}\displaystyle\sum_{n=0}^{\infty}\displaystyle\frac{q^{n+1/2}}{1-q^{2n+1}}\sin\left[(2n+1)\displaystyle\frac{\pi u}{2K}\right]\quad(|q|<1).$ (5.45)
Applying $q\equiv q\left(x\right)=e^{-y}$, then
${\rm sn}(u)$ $=\displaystyle\frac{2\pi}{\mathrm{K}k}\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{q^{n+\tfrac{1}{2}}\left(2n+1\right)^{2j+1}}{1-q^{2n+1}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j+1}}{\left(2j+1\right)!}}$
$=\displaystyle\frac{\pi}{\mathrm{K}k}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j+1}}{\left(2j+1\right)!}}S_{2j+1,1}^{\prime}\left(y\right).$ (5.46)
Comparing the coefficients of $u^{2m+1}$ in (5.44) and (5), we deduce
$S_{2n+1,1}^{\prime}\left(y\right)=\displaystyle\sum_{j=1}^{\infty}{\displaystyle\frac{\left(2j-1\right)^{2n+1}}{\sinh\left(\displaystyle\frac{2j-1}{2}y\right)}}=z^{2n+2}\sqrt{x}\displaystyle\frac{\mathrm{q}_{2n+1}\left(x\right)}{2}.$ (5.47)
Differentiating (5.47) yields
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}S_{2p-1,1}^{\prime}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{\left(2n-1\right)^{2p-1}}{\sinh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\text{DS}^{\prime}_{2p,2}\left(y\right).$
Using the relation $dx/dy=-x(1-x)z^{2}$ (see [5, P. 120, Entry. 9(i)]), we obtain
$\text{DS}^{\prime}_{2p,2}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}S_{2p-1,1}^{\prime}\left(y\right).$ (5.48)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we have
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0.$
Hence, the (5.47) and (5.48) can be rewritten as
$S_{2p-1,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)=\left(\displaystyle\frac{yz}{\pi}\right)^{2p}\displaystyle\frac{\mathrm{q}_{2p-1}\left(1-x\right)}{2}\sqrt{1-x}$ (5.49)
and
$\text{DS}^{\prime}_{2p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}S_{2p-1,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}S_{2p-1,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.50)
Finally, combining (5.43), (5.47), (5.49) and (5) yields the desired result with an elementary calculation. ∎
By Mathematica we can compute the following easily using the formulas in Theorem 5.1.
Corollary 5.2.
([23, Cor. 1]) We have
$\bar{C}_{2,2}(y)$ $=\displaystyle\frac{1}{8}\sqrt{1-x}xz^{3}[4(x-1)z^{\prime}+z],$
$\bar{C}_{4,2}(y)$ $=\displaystyle\frac{1}{32}\sqrt{1-x}xz^{5}[8(x^{2}-3x+2)z^{\prime}+(3x-4)z],$
$\bar{C}_{6,2}(y)$ $=\displaystyle\frac{1}{128}\sqrt{1-x}xz^{7}[12(x^{3}-17x^{2}+32x-16)z^{\prime}+(5x^{2}-52x+48)z],$
$\bar{C}_{8,2}(y)$ $=\displaystyle\frac{1}{512}\sqrt{1-x}xz^{9}\left\{16(x^{4}-139x^{3}+546x^{2}-680x+272)z^{\prime}\right.$
$\quad\quad\quad\quad\left.+(7x^{3}-696x^{2}+1776x-1088)z\right\},$
$\bar{C}_{10,2}(y)$ $=\displaystyle\frac{1}{2048}\sqrt{1-x}xz^{11}\left\{20(x^{5}-1233x^{4}+10400x^{3}-25040x^{2}+23808x-7936)z^{\prime}\right.$
$\quad\quad\quad\quad\left.+(9x^{4}-8632x^{3}+53232x^{2}-84288x+39680)z\right\}.$
Theorem 5.3.
Let $p\geq 2$ be a positive integer. We have
$\bar{S}_{2p,2}\left(y\right)$ $=\left(-1\right)^{p-1}x\left(1-x\right)\displaystyle\frac{z^{2p+2}}{2^{2p}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}$
$\quad+2p\left(-1\right)^{p-1}\left(1-x\right)^{2}x\displaystyle\frac{z^{2p+1}z^{\prime}}{2^{2p}}\mathrm{Q}_{2p-2}\left(1-x\right),$ (5.51)
where $\text{Q}_{2p-2}\left(x\right)$ appears in the coefficients of the Maclaurin series of $\mathrm{sn}^{2}\left(u\right)$.
Proof.
According to Lemma 3.4, we have
$\bar{S}_{2p,2}\left(y\right)=\displaystyle\frac{2p\left(-1\right)^{p-1}\pi^{2p}}{y^{2p+1}}S_{2p-1,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)+\left(-1\right)^{p}\displaystyle\frac{\pi^{2p+2}}{y^{2p+2}}\text{DS}_{2p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.52)
From [9], we have
$\mathrm{sn}^{2}u$ $=\left(1-\displaystyle\frac{E}{K}\right)\displaystyle\frac{1}{k^{2}}-\displaystyle\frac{\pi^{2}}{K^{2}k^{2}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{n\cos(n\pi u/K)}{\sinh(ny)}$
$=\left(1-\displaystyle\frac{E}{K}\right)\displaystyle\frac{1}{k^{2}}-\displaystyle\frac{\pi^{2}}{K^{2}k^{2}}\displaystyle\sum_{j=0}^{\infty}\displaystyle\frac{(-1)^{j}(\pi u/K)^{2j}}{(2j)!}S_{2j+1,1}(y),$
where $E$ is the complete elliptic integrals of the second kinds, see Section 2. Applying (5.44), we deduce the power series expansion of the function $\mathrm{sn}^{2}(u)$
$\mathrm{sn}^{2}u$ $=\left(\displaystyle\sum_{n=0}^{\infty}{\mathrm{q}_{2n+1}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n+1}}{\left(2n+1\right)!}}\right)^{2}$
$=\displaystyle\sum_{n=0}^{\infty}{\displaystyle\sum_{j=0}^{n}{\left(\begin{array}[]{c}2n+2\\ 2j+1\\ \end{array}\right)\mathrm{q}_{2j+1}\left(x\right)\mathrm{q}_{2n-2j+1}\left(x\right)}}\displaystyle\frac{\left(-1\right)^{n}u^{2n+2}}{\left(2n+2\right)!}=\displaystyle\sum_{n=0}^{\infty}{\mathrm{Q}_{2n+2}\left(x\right)}\displaystyle\frac{\left(-1\right)^{n}u^{2n+2}}{\left(2n+2\right)!}.$
Comparing the coefficients of $u^{2p-2}$ yields
$S_{2p-1,1}\left(y\right)=\displaystyle\frac{z^{2p}x}{2^{2p}}\mathrm{Q}_{2p-2}\left(x\right)\quad(p=2,3,4,\ldots).$ (5.53)
Differentiating (5.53) gives
$\text{DS}_{2p,2}\left(y\right)$ $=-\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}S_{2p-1,1}\left(y\right)=-\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\left[\displaystyle\frac{z^{2p}x}{2^{2p}}\mathrm{Q}_{2p-2}\left(x\right)\right]$
$=-\displaystyle\frac{\mathrm{d}x}{\mathrm{d}y}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\displaystyle\frac{z^{2p}x}{2^{2p}}\mathrm{Q}_{2p-2}\left(x\right)\right]=x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\displaystyle\frac{z^{2p}x}{2^{2p}}\mathrm{Q}_{2p-2}\left(x\right)\right]$
$=x\left(1-x\right)\displaystyle\frac{z^{2p+2}}{2^{2p}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[x\mathrm{Q}_{2p-2}\left(x\right)\right]+2px^{2}\left(1-x\right)\displaystyle\frac{z^{2p+1}z^{\prime}}{2^{2p}}\mathrm{Q}_{2p-2}\left(x\right).$ (5.54)
Applying the transformation
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0,$ (5.55)
by direct calculations, we obtain
$S_{2p-1,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=\displaystyle\frac{y^{2p}z^{2p}\left(1-x\right)}{2^{2p}\pi^{2p}}\mathrm{Q}_{2p-2}\left(1-x\right)$ (5.56)
and
$\text{DS}_{2p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-x\left(1-x\right)\displaystyle\frac{y^{2p+2}z^{2p+2}}{2^{2p}\pi^{2p+2}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}$
$\quad+2p\left(1-x\right)^{2}x\displaystyle\frac{y^{2p+1}z^{2p+1}}{2^{2p}\pi^{2p+2}}\left[\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right]\mathrm{Q}_{2p-2}\left(1-x\right)$
$=-x\left(1-x\right)\displaystyle\frac{y^{2p+2}z^{2p+2}}{2^{2p}\pi^{2p+2}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}$
$\quad+2p\left(1-x\right)\displaystyle\frac{y^{2p+1}z^{2p}}{2^{2p}\pi^{2p+2}}\mathrm{Q}_{2p-2}\left(1-x\right)+2p\left(1-x\right)^{2}x\displaystyle\frac{y^{2p+2}z^{2p+1}z^{\prime}}{2^{2p}\pi^{2p+2}}\mathrm{Q}_{2p-2}\left(1-x\right).$ (5.57)
Finally, substituting (5.56) and (5) into (5.52) yields the desired evaluation with a direct calculation. ∎
Setting $p=2,3,4,5$ in (5.3) with the help of Mathematica, we get the following evaluations.
Corollary 5.4.
([23, Cor.1]) We have
$\bar{S}_{4,2}(y)$ $=\displaystyle\frac{1}{8}x(1-x)z^{5}[z-4(1-x)z^{\prime}],$
$\bar{S}_{6,2}(y)$ $=\displaystyle\frac{1}{8}x(1-x)z^{7}[6(1-x)(2-x)z^{\prime}+(2x-3)z],$
$\bar{S}_{8,2}(y)$ $=\displaystyle\frac{1}{8}x(1-x)z^{9}[4(2x^{3}-19x^{2}+34x-17)z^{\prime}+(3x^{2}-19x+17)z],$
$\bar{S}_{10,2}(y)$ $=\displaystyle\frac{1}{8}x(1-x)z^{11}\left\{\begin{array}[]{c}10(x^{4}-34x^{3}+126x^{2}-155x+62)z^{\prime}\\ +(4x^{3}-102x^{2}+252x-155)z\\ \end{array}\right\}.$
From [23, Cor.1], we also obtain
$\bar{S}_{2,2}(y)=\displaystyle\frac{1}{8}x(1-x)z^{2}\left\{4x(1-x)\left(z^{\prime}\right)^{2}+4\left(1-x\right)z^{\prime}z-z^{2}\right\}.$
6 Evaluations of Mixed Berndt-type Integrals
Assuming that $x_{0},y_{0}$ and $z_{0}$ are all known. For $\Omega\left(x_{0},\exp\left(-y_{0}\right),z_{0},z_{0}^{\prime}\right)=0$, considering the transformation (3.18)
$\Omega\left(\left(\displaystyle\frac{1-\sqrt{1-x}}{1+\sqrt{1-x}}\right)^{2},\exp\left(-2y\right),\displaystyle\frac{z\left(1+\sqrt{1-x}\right)}{2},\displaystyle\frac{\left(1+\sqrt{1-x}\right)^{3}}{4\left(1-\sqrt{1-x}\right)}\left(\left(\sqrt{1-x}+1-x\right)z^{\prime}-\displaystyle\frac{z}{2}\right)\right)=0,$
by an elementary calculation, we can deduce its inverse transformation
$\begin{cases}x=\displaystyle\frac{4\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}},\\ y=\displaystyle\frac{y_{0}}{2},\\ z=z_{0}\left(1+\sqrt{x_{0}}\right),\\ z^{\prime}=\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}.\end{cases}$ (6.58)
Theorem 6.1.
Let $m\in\mathbb{N}$ and $\Gamma=\Gamma(1/4)$. Then
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=c_{1,m}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+c_{2,m}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$ (6.59)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sin x+\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=d_{1,m}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+d_{2,m}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}},$ (6.60)
where numbers $c_{1,m},c_{2,m},d_{1,m},d_{2,m}\in\mathbb{Q}$.
Proof.
According to Lemma LABEL:sn_coefficient_symmetry, we know that $\text{q}_{2p-1}(x)=\displaystyle\sum^{p-1}_{j=0}a_{j}x^{j}$. Applying Theorem 5.1 and transformation (6.58), one obtains
$\bar{C}_{2p,2}\left(y\right)=\bar{C}_{2p,2}\left(\displaystyle\frac{y_{0}}{2}\right)$
$=\left(-1\right)^{p}\displaystyle\frac{px(1-x)}{2^{2p-1}}\mathrm{q}_{2p-1}\left(1-x\right)\sqrt{1-x}z^{2p+1}z^{\prime}+\left(-1\right)^{p}\displaystyle\frac{x(1-x)}{2^{2p}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\mathrm{q}_{2p-1}\left(1-x\right)\sqrt{1-x}\right]z^{2p+2}$
$=\left(-1\right)^{p}\displaystyle\frac{p\sqrt{x_{0}}\left(1-\sqrt{x_{0}}\right)^{3}}{2^{2p-3}\left(1+\sqrt{x_{0}}\right)^{4-2p}}\mathrm{q}_{2p-1}\left(\displaystyle\frac{\left(1-\sqrt{x_{0}}\right)^{2}}{\left(1+\sqrt{x_{0}}\right)^{2}}\right)z_{0}^{2p+1}\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}$
$\quad-\displaystyle\frac{\left(-1\right)^{p}}{2^{2p-2}}\displaystyle\frac{\sqrt{x_{0}}\left(1-\sqrt{x_{0}}\right)^{2}}{\left(1+\sqrt{x_{0}}\right)^{4}}\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left(j+\displaystyle\frac{1}{2}\right)\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2j-1}\left(1+\sqrt{x_{0}}\right)^{2p+2}z_{0}^{2p+2}$
$=\left(-1\right)^{p}\displaystyle\frac{px_{0}\left(1-x_{0}\right)^{2}}{2^{2p-2}}\left(1+\sqrt{x_{0}}\right)^{2p-2}\mathrm{q}_{2p-1}\left(\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2}\right)z_{0}^{2p+1}z_{0}^{\prime}$
$\quad-\displaystyle\frac{\left(-1\right)^{p}}{2^{2p-2}}(1-x_{0})^{2}\sqrt{x_{0}}\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left(j+\displaystyle\frac{1}{2}\right)\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2j-1}\left(1+\sqrt{x_{0}}\right)^{2p-4}z_{0}^{2p+2}$
$\quad+\displaystyle\frac{\left(-1\right)^{p}}{2^{2p-2}}(1-x_{0})^{2}\displaystyle\frac{p}{2}\sqrt{x_{0}}\mathrm{q}_{2p-1}\left(\displaystyle\frac{\left(1-\sqrt{x_{0}}\right)^{2}}{\left(1+\sqrt{x_{0}}\right)^{2}}\right)\left(1+\sqrt{x_{0}}\right)^{2p-3}z_{0}^{2p+2}.$ (6.61)
If letting $x_{0}=1/2$ in (6.61), then $y_{0}=\pi$ by (2.14). And the coefficients before $z_{0}^{2p+2},z_{0}^{2p+1}z_{0}^{\prime}$ are all rational numbers. when $x_{0}=1/2$, the coefficient of $z_{0}^{2p+1}z_{0}^{\prime}$ is
$\left(1+\displaystyle\frac{\sqrt{2}}{2}\right)^{2p-2}\displaystyle\frac{\left(-1\right)^{p}p}{2^{2p+1}}\mathrm{q}_{2p-1}\left(\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2}\right)$
$=\displaystyle\frac{\left(-1\right)^{p}p}{2^{3p}}\left(\sqrt{2}+1\right)^{2p-2}\displaystyle\sum_{j=0}^{p-1}{a_{j}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j}}.$ (6.62)
In fact, Lemma LABEL:sn_coefficient_symmetry indicates that $a_{j}=a_{p-1-j}$, for $j=0,\ldots,p-1$, Thus,
$\left(\sqrt{2}+1\right)^{2p-2}$$\left\{a_{j}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j}+a_{p-1-j}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2\left(p-1-j\right)}\right\}$
$=a_{j}\left\{\left(\sqrt{2}-1\right)^{4j+2p-2}+\left(\sqrt{2}+1\right)^{2p-2}\left(\sqrt{2}-1\right)^{4\left(p-1-j\right)}\right\}$
$=a_{j}\left\{\left(\sqrt{2}-1\right)^{4j+2p-2}+\left(\sqrt{2}+1\right)^{4j+2p-2}\right\}$
$=a_{j}\displaystyle\sum_{l=0}^{4j+2p-2}{\left(\begin{array}[]{c}4j+2p-2\\ l\\ \end{array}\right)}\left[\left(-1\right)^{l}+1\right]2^{2j+p-1-\tfrac{l}{2}}$ (6.65)
$=2a_{j}\displaystyle\sum_{m=0}^{2j+p-1}{\left(\begin{array}[]{c}4j+2p-2\\ 2m\\ \end{array}\right)}2^{2j+p-1-m}\in\mathbb{Q}\quad(j=1,\ldots,[(p-1)/2]).$ (6.68)
Further, for $j=0,\ldots,p-1$, we obtain
$\displaystyle\frac{\left(-1\right)^{p}p}{2^{3p}}\left(\sqrt{2}+1\right)^{2p-2}\displaystyle\sum_{j=0}^{p-1}{a_{j}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j}}\in\mathbb{Q}.$ (6.69)
Similarly, when $x_{0}=1/2$, the coefficient of $z_{0}^{2p+2}$ is
$\quad\displaystyle\frac{\left(-1\right)^{p}}{2^{2p+1}}\sqrt{2}\left\{\begin{array}[]{c}\displaystyle\frac{p}{2}\mathrm{q}_{2p-1}\left(\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2}\right)\left(1+\displaystyle\frac{\sqrt{2}}{2}\right)^{2p-3}\\ -\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left(j+\displaystyle\frac{1}{2}\right)\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j-1}\left(1+\displaystyle\frac{\sqrt{2}}{2}\right)^{2p-4}\\ \end{array}\right\}$ (6.72)
$\quad\quad=\displaystyle\frac{\left(-1\right)^{p}}{2^{3p+1}}\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left\{\begin{array}[]{c}\sqrt{2}p\left(1+\sqrt{2}\right)\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j}\\ -\left(4j+2\right)\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j-1}\\ \end{array}\right\}$ (6.75)
$\quad\quad=\displaystyle\frac{\left(-1\right)^{p}}{2^{3p+1}}\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left\{\left(2+\sqrt{2}\right)p-\left(4j+2\right)\right\}\left(\sqrt{2}-1\right)^{4j-2}.$ (6.76)
According to $a_{j}=a_{p-1-j},j=0,\ldots,p-1$, we have
$\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}$$\left\{\begin{array}[]{c}a_{j}\left\{\left(2+\sqrt{2}\right)p-2\left(2j+1\right)\right\}\left(\sqrt{2}-1\right)^{4j-2}+\\ a_{p-1-j}\left\{\left(2+\sqrt{2}\right)p-2\left(2p-2j-1\right)\right\}\left(\sqrt{2}-1\right)^{4p-4j-6}\\ \end{array}\right\}$ (6.79)
$=a_{j}\left\{\begin{array}[]{c}\left\{\left(2\sqrt{2}+4\right)p-2\sqrt{2}\left(2j+1\right)\right\}\left(\sqrt{2}+1\right)^{2p-4j-2}+\\ \left\{\left(2\sqrt{2}+4\right)p-2\sqrt{2}\left(2p-2j-1\right)\right\}\left(\sqrt{2}-1\right)^{2p-4j-2}\\ \end{array}\right\}$ (6.82)
$=a_{j}\left\{\begin{array}[]{c}2\sqrt{2}\left(2j+1\right)\left\{\left(\sqrt{2}-1\right)^{2p-4j-2}-\left(\sqrt{2}+1\right)^{2p-4j-2}\right\}\\ 2\sqrt{2}p\left\{\left(\sqrt{2}-1\right)^{2p-4j-1}+\left(\sqrt{2}+1\right)^{2p-4j-1}\right\}\\ \end{array}\right\}$ (6.85)
$=4a_{j}\left\{\begin{array}[]{c}p\displaystyle\sum_{l=0}^{p-2j-1}{\left(\begin{array}[]{c}2p-4j-1\\ 2l\\ \end{array}\right)2^{p-2j-l}}\\ -\left(2j+1\right)\displaystyle\sum_{l=0}^{p-2j-2}{\left(\begin{array}[]{c}2p-4j-2\\ 2l+1\\ \end{array}\right)2^{p-2j-1-l}}\\ \end{array}\right\}\in\mathbb{Q},$ (6.92)
where $j=1,\ldots,[(p-1)/2]$.
Combining (6.69) and (6.92) gives
$\bar{C}_{2p,2}\left(\displaystyle\frac{\pi}{2}\right)\in\mathbb{Q}z_{0}^{2p+1}z_{0}^{\prime}+\mathbb{Q}z_{0}^{2p+2}.$ (6.93)
When $x_{0}=1/2$ then $y_{0}=\pi,\ z_{0}=\displaystyle\frac{\Gamma^{2}}{2\pi^{3/2}}$ and $z_{0}^{\prime}=\displaystyle\frac{4\pi^{1/2}}{\Gamma^{2}}$ by (2.14), then substitute it into (6.93) yields
$\bar{C}_{2p,2}\left(\displaystyle\frac{\pi}{2}\right)\in\mathbb{Q}\displaystyle\frac{\Gamma^{4p}}{\pi^{3p+1}}+\mathbb{Q}\displaystyle\frac{\Gamma^{4p+4}}{\pi^{3p+3}}.$ (6.94)
In Theorem 4.1, letting $a=4m-2$ and $4m$, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\left(-1\right)^{m-1}\displaystyle\frac{\pi^{4m-1}}{4^{m}}\bar{C}_{4m-2,2}\left(\displaystyle\frac{\pi}{2}\right),$ (6.95)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\left(-1\right)^{m-1}\displaystyle\frac{\pi^{4m+1}}{2\cdot 4^{m}}\bar{C}_{4m,2}\left(\displaystyle\frac{\pi}{2}\right).$ (6.96)
Finally, substituting (6.94) into (6.95) and (6.96), respectively, we prove the theorem. ∎
By Mathematica we can compute the following easily using the formulas in Theorem 6.1:
Example 6.2.
Let $\Gamma=\Gamma(1/4)$. We have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{\Gamma^{8}}{2^{9}\pi^{3}}-\displaystyle\frac{\Gamma^{4}}{2^{6}\pi},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{6}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{9\Gamma^{12}}{16384\pi^{3}}-\displaystyle\frac{3\Gamma^{16}}{131072\pi^{5}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{10}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{459\Gamma^{24}}{33554432\pi^{7}}-\displaystyle\frac{945\Gamma^{20}}{4194304\pi^{5}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{3\Gamma^{8}}{1024\pi^{2}}-\displaystyle\frac{\Gamma^{12}}{8192\pi^{4}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{8}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{33\Gamma^{20}}{2097152\pi^{6}}-\displaystyle\frac{63\Gamma^{16}}{262144\pi^{4}}.$
Theorem 6.3.
Let $m\in\mathbb{N}$ and $\Gamma=\Gamma(1/4)$. Then
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=e_{1,m}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+e_{2,m}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$ (6.97)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=f_{1,m}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+f_{2,m}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}}\quad(m\geq 2).$ (6.98)
where the numbers $e_{1,m},e_{2,m},f_{1,m},f_{2,m}\in\mathbb{Q}$.
Proof.
Applying Theorem 5.3 and (6.58), by a direct calculation, we deduce
$\bar{S}_{2p,2}\left(y\right)=\bar{S}_{2p,2}\left(\displaystyle\frac{y_{0}}{2}\right)$ (6.99)
$=\displaystyle\frac{\left(-1\right)^{p-1}x\left(1-x\right)}{2^{2p}}\left[\begin{array}[]{c}z^{2p+2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}\\ +2p\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)z^{2p+1}z^{\prime}\\ \end{array}\right]$ (6.102)
$=\displaystyle\frac{\left(-1\right)^{p-1}\displaystyle\frac{4\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}}{2^{2p}}{z_{0}}^{2p+1}\left(1+\sqrt{x_{0}}\right)^{2p+1}$
×$\quad\quad\times\left[\begin{array}[]{c}z_{0}\left(1+\sqrt{x_{0}}\right)\displaystyle\lim_{x=4\sqrt{x_{0}}/(1+x_{0}+2\sqrt{x_{0}})}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}\\ +2p\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\mathrm{Q}_{2p-2}\left(\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\right)\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}\\ \end{array}\right]$ (6.105)
$=\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{4}{z_{0}}^{2p+2}}{2^{2p-3}}p\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-9}\mathrm{Q}_{2p-2}\left(\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2}\right)$
×$\quad\quad\times\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}+\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}{z_{0}}^{2p+2}}{2^{2p-2}}\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-4}$
×$\quad\quad\times\displaystyle\lim_{x=4\sqrt{x_{0}}/(1+x_{0}+2\sqrt{x_{0}})}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}.$ (6.106)
When $x_{0}=1/2$ in (6.106), we prove that the coefficients before $z_{0}^{2p+2},z_{0}^{2p+1}z_{0}^{\prime}$ are all rational numbers. The coefficient before $z_{0}^{\prime}{z_{0}}^{2p+1}$ is
$\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}}{2^{2p-1}}$ $\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-5}\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}2p\displaystyle\frac{1+x-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\mathrm{Q}_{2p-2}\left(\displaystyle\frac{1+x-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\right)$
$=\displaystyle\frac{\left(-1\right)^{p-1}p\left(1-x_{0}\right)^{3}x_{0}}{2^{2p-2}}\left(1+\sqrt{x_{0}}\right)^{2p-4}\mathrm{Q}_{2p-2}\left(\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2}\right)$
$=\displaystyle\frac{\left(-1\right)^{p-1}p\left(1-x_{0}\right)^{3}x_{0}}{2^{2p-2}}\left(1+\sqrt{x_{0}}\right)^{2p-4}\displaystyle\sum_{j=0}^{p-2}{b_{j}\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2j}}.$ (6.107)
Applying $a_{j}=a_{p-1-j}$ and $\mathrm{Q}_{2p-2}\left(x\right)=\displaystyle\sum_{j=0}^{p-2}{\left(\begin{array}[]{c}2p-2\\ 2j+1\\ \end{array}\right)\mathrm{q}_{2j+1}\left(x\right)\mathrm{q}_{2p-2j-3}\left(x\right)}$, we have
$\mathrm{Q}_{2p-2}\left(x\right)=\displaystyle\sum_{j=0}^{p-2}{b_{j}x^{j}}\quad(b_{j}\in\mathbb{Q}),$ (6.108)
wich can reaily reduce that $b_{j}=b_{p-j-2}\ (j=0,\ldots,p-2)$. Thus,
$\left(\sqrt{2}+1\right)^{2p-4}$$\left\{b_{j}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j}+b_{p-j-2}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2\left(p-j-2\right)}\right\}$
$=b_{j}\left\{\left(\sqrt{2}-1\right)^{4j-2p+4}+\left(\sqrt{2}+1\right)^{2p-4}\left(\sqrt{2}-1\right)^{4\left(p-j-2\right)}\right\}$
$=b_{j}\left\{\left(\sqrt{2}-1\right)^{4j-2p+4}+\left(\sqrt{2}+1\right)^{4j-2p+4}\right\}$
$=b_{j}\displaystyle\sum_{l=0}^{4j-2p+4}{\left(\begin{array}[]{c}4j-2p+4\\ l\\ \end{array}\right)}\left[\left(-1\right)^{l}+1\right]2^{2j-p+2-\tfrac{l}{2}}$ (6.111)
$=2b_{j}\displaystyle\sum_{m=0}^{2j-p+2}{\left(\begin{array}[]{c}4j-2p+4\\ 2m\\ \end{array}\right)}2^{2j-p+2-m}\in\mathbb{Q}\quad(j=0,\ldots,p-2).$ (6.114)
Hence, we obtain
$\displaystyle\frac{\left(-1\right)^{p-1}p}{2^{3p}}\left(\sqrt{2}+1\right)^{2p-4}\displaystyle\sum_{j=0}^{p-2}{b_{j}\left(\displaystyle\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)^{2j}}\in\mathbb{Q}.$ (6.115)
Similarly, the coefficient of ${z_{0}}^{2p+2}$ is
$\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}}{2^{2p-2}}\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-5}\left[\begin{array}[]{c}\left(1+\sqrt{x_{0}}\right)\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}\\ +\displaystyle\frac{p}{2}\left(1-x_{0}\right)\mathrm{Q}_{2p-2}\left(\left(\displaystyle\frac{1-\sqrt{x_{0}}}{1+\sqrt{x_{0}}}\right)^{2}\right)\\ \end{array}\right].$ (6.118)
Applying (6.108) to (6.118), we have
$\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}}{2^{2p-2}}\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-5}\displaystyle\sum_{j=0}^{p-2}{b_{j}}\left[\displaystyle\frac{p}{2}\left(1-x_{0}\right)-\left(1+\sqrt{x_{0}}\right)\left(j+1\right)\right]\left(\sqrt{2}-1\right)^{4j}$
$=\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}}{2^{2p-2}}\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-4}\displaystyle\sum_{j=0}^{p-2}{b_{j}}\left[\displaystyle\frac{p}{2}\left(1-\sqrt{x_{0}}\right)-\left(j+1\right)\right]\left(\sqrt{2}-1\right)^{4j}$
$=\displaystyle\frac{\left(-1\right)^{p-1}}{2^{3p-2}}\left(\sqrt{2}+1\right)^{2p-4}\displaystyle\sum_{j=0}^{p-2}{b_{j}}\left[\displaystyle\frac{p}{4}\left(\sqrt{2}-1\right)-\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\right]\left(\sqrt{2}-1\right)^{4j}.$ (6.119)
Noting the fact that
$\left(\sqrt{2}+1\right)^{2p-4}\left\{\begin{array}[]{c}\left[\displaystyle\frac{p}{4}\left(\sqrt{2}-1\right)-\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\right]\left(\sqrt{2}-1\right)^{4j}+\\ \left[\displaystyle\frac{p}{4}\left(\sqrt{2}-1\right)-\displaystyle\frac{\sqrt{2}}{2}\left(p-j-1\right)\right]\left(\sqrt{2}-1\right)^{4\left(p-j-2\right)}\\ \end{array}\right\}$ (6.122)
$=\left[\displaystyle\frac{p}{4}\left(\sqrt{2}-1\right)-\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\right]\left(\sqrt{2}+1\right)^{2p-4j-4}+\left[-\displaystyle\frac{p}{4}\left(\sqrt{2}+1\right)+\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\right]\left(\sqrt{2}-1\right)^{2p-4j-4}$
$=\displaystyle\frac{p}{4}\left[\left(\sqrt{2}+1\right)^{2p-4j-5}-\left(\sqrt{2}-1\right)^{2p-4j-5}\right]+\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\left[\left(\sqrt{2}-1\right)^{2p-4j-4}-\left(\sqrt{2}+1\right)^{2p-4j-4}\right],$
$=\displaystyle\frac{p}{4}\displaystyle\sum_{l=0}^{2p-4j-5}{\left(\begin{array}[]{c}2p-4j-5\\ l\\ \end{array}\right)\left[1-\left(-1\right)^{l}\right]2^{p-2j-\tfrac{l+5}{2}}}+\displaystyle\frac{j+1}{2}\displaystyle\sum_{l=0}^{2p-4j-4}{\left(\begin{array}[]{c}2p-4j-4\\ l\\ \end{array}\right)\left[\left(-1\right)^{l}-1\right]2^{p-2j-\tfrac{l+3}{2}}}$ (6.127)
$=\displaystyle\frac{p}{2}\displaystyle\sum_{l=0}^{p-2j-3}{\left(\begin{array}[]{c}2p-4j-5\\ 2l+1\\ \end{array}\right)2^{p-2j-3-l}}-\left(j+1\right)\displaystyle\sum_{l=0}^{p-2j-2}{\left(\begin{array}[]{c}2p-4j-4\\ 2l+1\\ \end{array}\right)2^{p-2j-2-l}}\in\mathbb{Q},$ (6.132)
where $j=0,\ldots,[(p-2)/2]$.
Combining (6.115) with (6.122), we derive that
$\bar{S}_{2p,2}\left(\displaystyle\frac{\pi}{2}\right)\in\mathbb{Q}z_{0}^{2p+1}z_{0}^{\prime}+\mathbb{Q}z_{0}^{2p+2}\quad(p\geq 2).$ (6.133)
Substituting $z_{0}=\displaystyle\frac{\Gamma^{2}}{2\pi^{3/2}}$ and $z_{0}^{\prime}=\displaystyle\frac{4\pi^{1/2}}{\Gamma^{2}}$ into (6.133) gives
$\bar{S}_{2p,2}\left(\displaystyle\frac{\pi}{2}\right)\in\mathbb{Q}\displaystyle\frac{\Gamma^{4p}}{\pi^{3p+1}}+\mathbb{Q}\displaystyle\frac{\Gamma^{4p+4}}{\pi^{3p+3}}.$ (6.134)
In Theorem 4.2, letting $a=4m$ and $24m-$ 2, we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=\left(-1\right)^{m-1}\displaystyle\frac{\pi^{4m+1}}{2\cdot 4^{m}}\bar{S}_{4m,2}\left(\displaystyle\frac{\pi}{2}\right),$ (6.135)
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=\left(-1\right)^{m-1}\displaystyle\frac{\pi^{4m-1}}{4^{m}}\bar{S}_{4m-2,2}\left(\displaystyle\frac{\pi}{2}\right).$ (6.136)
Finally, substituting (6.134) into (6.135) and (6.136) yields the desired result. ∎
Remark 6.4.
In particular, by using the methods of the proof in Theorem 6.3 and the explicit expression of $\bar{S}_{2,2}\left(\displaystyle\frac{\pi}{2}\right)$, we can find that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=\displaystyle\frac{\Gamma^{4}}{2^{6}\pi}-\displaystyle\frac{\pi}{8}.$ (6.137)
By Mathematica we can compute the following easily using the formulas in Theorem 6.3:
Example 6.5.
Let $\Gamma=\Gamma(1/4)$. Then
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=\displaystyle\frac{\Gamma^{12}}{2^{13}\pi^{4}}-\displaystyle\frac{\Gamma^{8}}{2^{10}\pi^{2}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{8}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=\displaystyle\frac{81\Gamma^{16}}{262144\pi^{4}}-\displaystyle\frac{33\Gamma^{20}}{2097152\pi^{6}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{6}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=\displaystyle\frac{5\Gamma^{16}}{131072\pi^{5}}-\displaystyle\frac{9\Gamma^{12}}{16384\pi^{3}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{10}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=\displaystyle\frac{945\Gamma^{20}}{4194304\pi^{5}}-\displaystyle\frac{405\Gamma^{24}}{33554432\pi^{7}}.$
Corollary 6.6.
For $m\in\mathbb{N}$, we have
$\pi\bar{S}_{4m,2}\left(\displaystyle\frac{\pi}{2}\right)$ $=4^{2m+1}\left[\pi\mathrm{DS}_{4m,2}\left(2\pi\right)-2mS_{4m-1,1}\left(2\pi\right)\right],$ (6.138)
$\pi\bar{S}_{4m-2,2}\left(\displaystyle\frac{\pi}{2}\right)$ $=4^{2m}\left[\left(2m-1\right)S_{4m-3,1}\left(2\pi\right)-\pi\mathrm{DS}_{4m-2,2}\left(2\pi\right)\right],$ (6.139)
$\pi\bar{C}_{4m-2,2}\left(\displaystyle\frac{\pi}{2}\right)$ $=4\left[\pi\mathrm{DS}_{4m-2,2}^{\prime}\left(2\pi\right)-\left(4m-2\right)S_{4m-3,1}^{\prime}\left(2\pi\right)\right],$ (6.140)
$\pi\bar{C}_{4m,2}\left(\displaystyle\frac{\pi}{2}\right)$ $=4\left[4mS_{4m-1,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2}^{\prime}\left(2\pi\right)\right].$ (6.141)
Proof.
The facts derived from Theorem 4.3 and Theorem 4.4 lead us to the conclusion that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left[\sinh x-\sin x\right]\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=\left(-4\right)^{m}\pi^{4m-2}\left[\pi\mathrm{DS}_{4m-2,2,1}\left(2\pi\right)-\left(2m-1\right)S_{4m-3,1}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left[\sinh x+\sin x\right]\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=2\left(-4\right)^{m}\pi^{4m}\left[2mS_{4m-1,1}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2,1}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\left(-1\right)^{m}}{2^{2m-2}}\pi^{4m-2}\left[\left(4m-2\right)S_{4m-3,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m-2,2,1}^{\prime}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\left(-1\right)^{m-1}}{2^{2m-1}}\pi^{4m}\left[4mS_{4m-1,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2,1}^{\prime}\left(2\pi\right)\right].$ (6.142)
Hence, combining (6.95),(6.96),(6.135),(6.136) and (6), the desired result is established. ∎
Finally, we conclude this section by the following conjecture which is supported by comparing the coefficients in Examples 6.2 and 6.5,
Conjecture 6.7.
For $m\geq 1$, let $c_{2,m},e_{2,m},d_{1,m},f_{1,m}$ be the numbers appearing in Theorems 6.1 and 6.3. Then the following relations hold
$c_{2,m}+e_{2,m}=0\quad\text{and}\quad d_{1,m}+f_{1,m}=0.$
7 Berndt-type Integrals via Barnes Multiple Zeta Functions
In this section, we evaluate mixed Berndt-type integrals using the generalized Barnes multiple zeta function. Bradshaw and Vignat [7] evaluated generalized Berndt-type integrals (1.1) using the (alternating) Barnes’ multiple zeta function. For example, they found that ([7, Prop. 2])
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}dx}{(\cos x-\cosh x)^{b}}=2^{b}\Gamma(a+1)\zeta_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 2b,b\geq 1)$ (7.143)
and
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}dx}{(\cos x+\cosh x)^{b}}=2^{b}\Gamma(a+1){\bar{\zeta}}_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 0,b\geq 1),$ (7.144)
[…]
$\zeta_{N}(s,\omega|{\bf a}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{1}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\Re(s)>N)$ (7.145)
and
${\bar{\zeta}}_{N}(s,\omega|{\bf a}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{(-1)^{n_{1}+\cdots+n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\Re(s)>N-1),$ (7.146)
respectively.
In particular, Z.P. Bradshaw and C. Vignat [7] gave the integral representation of Barnes multiple zeta function, which appears as [15, Eq. (3.2)].
Proposition 7.1.
Let $\Re(s)>N,\Re(w)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1-e^{-a_{j}u})^{-1}du.$
The alternating version is
$\widetilde{\zeta}_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1+e^{-a_{j}u})^{-1}du.$
Here we define the generalized Barnes multiple zeta function by
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N}):=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\displaystyle\frac{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}}\quad(\Re(\omega)>\Re(s)>N),\quad$ (7.147)
where $\boldsymbol{\sigma}_{N}=\left(\sigma_{1},\ldots,\sigma_{N}\right)\in\left\{\pm 1\right\}^{N}$, and find its an integral representation:
Proposition 7.2.
Let $\Re(s)>N,\Re(w)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})=\displaystyle\frac{1}{\Gamma\left(s\right)}\displaystyle\int_{0}^{\infty}{u^{s-1}}e^{-wu}\prod_{j=1}^{N}{(}1-\sigma_{j}e^{-a_{j}u})^{-1}\mathrm{d}u,$
where $\boldsymbol{\sigma}_{N}=\left(\sigma_{1},\ldots,\sigma_{N}\right)\in\left\{\pm 1\right\}^{N}$.
Proof.
Since $\Re(s)>N$, we can write the gamma function in its integral representation, and the series representation of $\zeta_{N}$ converges absolutely for $\Re(w)>0$ and $\Re(a_{j})>0$. We have
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})\Gamma\left(s\right)$ $=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\displaystyle\frac{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\displaystyle\int_{0}^{\infty}{x^{s-1}e^{-x}\mathrm{d}x}}\quad$
$=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}x^{s-1}e^{-x}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}}\mathrm{d}x}$
$=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}\displaystyle\int_{0}^{\infty}{u^{s-1}e^{-(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})u}}\mathrm{d}u}.$
The sum can be moved inside by the dominated convergence theorem, namely
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})\Gamma\left(s\right)$ $=\displaystyle\int_{0}^{\infty}{u^{s-1}}e^{-\omega u}\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\left(\sigma_{1}e^{-a_{1}u}\right)^{n_{1}}\cdots\left(\sigma_{N}e^{-a_{N}u}\right)^{n_{N}}\mathrm{d}u}$
$=\displaystyle\int_{0}^{\infty}{u^{s-1}}e^{-\omega u}\prod_{j=1}^{N}{\left(1-\sigma_{j}e^{-a_{J}u}\right)^{-1}}\mathrm{d}u.$
Noting that $\Re(a_{j})>0$, the geometric series formula ensures that the last equality holds. ∎
Theorem 7.3.
For positive integer $m$, we have
$\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}\zeta_{4}\left(4m+1,3+i^{j}|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$
$\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right)$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}},$
$\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4})$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$
$\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4}\right)}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}}\quad(m\geq 2),$
where $\mathbf{a}_{4}=\left(2-2i,2+2i,1-i,1+i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$.
Proof.
Noting the facts that
$\displaystyle\frac{\left(\sinh x\pm\sin x\right)}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}=\displaystyle\frac{e^{x}-e^{-x}\pm e^{ix}\mp e^{-ix}}{4\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)},$
$\displaystyle\frac{\left(\sinh x\pm\sin x\right)}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}=\displaystyle\frac{e^{x}-e^{-x}\pm e^{ix}\mp e^{-ix}}{4\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)},$
From the proof of[7, Prop. 2], we have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{s-1}e^{-\omega x}}{\prod_{i=1}^{M}{\sinh\left(a_{1}x\right)}\prod_{j=1}^{N}{\cosh\left(b_{j}x\right)}}}\mathrm{d}x$ $=\displaystyle\sum_{\tiny\begin{array}[]{c}n_{1},\cdots,n_{M}\geq 0\\ k_{1},\cdots,k_{N}\geq 0\\ \end{array}}{\displaystyle\frac{2^{M+N}\Gamma\left(s\right)\left(-1\right)^{k_{1}+\cdots+k_{N}}}{\left(w+\displaystyle\sum_{i=1}^{M}{a_{i}+\displaystyle\sum_{i=1}^{N}{b_{i}+}2\left(\displaystyle\sum_{i=1}^{M}{a_{i}n_{i}}+\displaystyle\sum_{i=1}^{N}{b_{i}k_{i}}\right)}\right)^{s}}}$
$=2^{M+N}\Gamma\left(s\right)\zeta_{M+N}\left(s,w+\displaystyle\sum_{i=1}^{M}{a_{i}+}\displaystyle\sum_{i=1}^{N}{b_{i}}|\boldsymbol{c}_{M+N},\boldsymbol{\sigma}_{M+N}\right),$
where $\boldsymbol{c}_{M+N}=\left(2a_{1},\cdots,2a_{M},2b_{1},\cdots,2b_{N}\right),\boldsymbol{\sigma}_{M+N}=\left(\left\{1\right\}^{M},\left\{-1\right\}^{N}\right).$ Thus, we deduce that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sin x+\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(e^{x}-e^{-x}+e^{ix}-e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=\displaystyle\frac{1}{4}\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}e^{-i^{j}x}\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right),$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(e^{x}-e^{-x}-e^{ix}+e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=\displaystyle\frac{1}{4}\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}e^{-i^{j}x}\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}}\zeta_{4}\left(4m+1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right),$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(e^{x}-e^{-x}+e^{ix}-e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4}),$
and
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(e^{x}-e^{-x}-e^{ix}+e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4}),$
where $\mathbf{a}_{4}=\left(2-2i,2+2i,1-i,1+i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$.
Finally, using Theorems 6.1 and 6.3 yield the desired evaluations. ∎
Letting $m=1,2$, by Mathematica we can compute the following easily using the formulas in Theorem 7.3:
Example 7.4.
Let $\Gamma=\Gamma(1/4)$. Then
$\zeta_{4}\left(5,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(5,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(5,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(5,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
$=\displaystyle\frac{\Gamma^{8}}{32768\pi^{2}}-\displaystyle\frac{\Gamma^{12}}{786432\pi^{4}},$
$\zeta_{4}\left(9,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(9,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(9,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(9,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
$=\displaystyle\frac{11\Gamma^{20}}{112742891520\pi^{6}}-\displaystyle\frac{\Gamma^{16}}{671088640\pi^{4}},$
$\zeta_{4}\left(3,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(3,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(3,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(3,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
$=\displaystyle\frac{\Gamma^{8}}{2^{12}\pi^{3}}-\displaystyle\frac{\Gamma^{4}}{2^{9}\pi},$
$\zeta_{4}\left(7,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(7,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(7,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(7,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
$=\displaystyle\frac{\Gamma^{12}}{5242880\pi^{3}}-\displaystyle\frac{\Gamma^{16}}{125829120\pi^{5}},$
and
$\zeta_{4}(5,2|\mathbf{a}_{4})+\zeta_{4}(5,3-i|\mathbf{a}_{4})-\zeta_{4}(5,3+i|\mathbf{a}_{4})-\zeta_{4}(5,4|\mathbf{a}_{4})=\displaystyle\frac{\Gamma^{12}}{786432\pi^{4}}-\displaystyle\frac{\Gamma^{8}}{98304\pi^{2}},$
$\zeta_{4}(9,2|\mathbf{a}_{4})+\zeta_{4}(9,3-i|\mathbf{a}_{4})-\zeta_{4}(9,3+i|\mathbf{a}_{4})-\zeta_{4}(9,4|\mathbf{a}_{4})=\displaystyle\frac{9\Gamma^{16}}{4697620480\pi^{4}}-\displaystyle\frac{11\Gamma^{20}}{112742891520\pi^{6}},$
$\zeta_{4}(7,3+i|\mathbf{a}_{4})+\zeta_{4}(7,2|\mathbf{a}_{4})-\zeta_{4}(7,3-i|\mathbf{a}_{4})-\zeta_{4}(7,4|\mathbf{a}_{4})=\displaystyle\frac{\Gamma^{16}}{75497472\pi^{5}}-\displaystyle\frac{\Gamma^{12}}{5242880\pi^{3}},$
where $\mathbf{a}_{4}=\left(2-2i,2+2i,1-i,1+i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$.
Conflict of Interest. The author declare that they have no conflict of interest.
Data Availability Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Use of AI Tools Declaration. The author declare they have not used Artificial Intelligence (AI) tools in the creation of this article.